Write the point slope form of the equation of the line that passes through the points (6,-9) and (7, 1).
Found 2 solutions by woodychuck0022, rothauserc:
Answer by woodychuck0022(1)

1. Find the slope using the formula: y1-y2/x1-x2.
For this problem that would be 1--9/7-6, which equals the slope, 10.
2. Use the point slope form using the equation: y-y1=m(x-1).
You would plug in what you know, and have your answer.
3. Answer is: y-1=10(x-7)
*For finding y1, y2, x1, and x2, use the given ordered pairs. For step 3, either ordered paid can be used.

Answer by rothauserc(4718)

slope = (y2 -y1) / (x2 - x1) = (1 + 9) / (7 - 6) = 10
:
point one point, say (6, -9), then we have
:
(y + 9) = 10(x - 6)
: